OpenStudy (anonymous):

SOLVE: 2(4c-4)-5=6c+1 (please i need to find the answer ASAP) :D

OpenStudy (anonymous):

c=7

OpenStudy (mathstudent55):

Start by distributing the 2 on left side: 2(4c) - 2(4) - 5 = 6c +1 8c - 8 - 5 = 6c + 1 Now collect like terms 8c - 13 = 6c + 1 Now add 13 to both sides and subtract 6c from both sides 2c = 14 Now divide both sides by 2 c = 7

OpenStudy (anonymous):

Thanks to both!!